# Towards Human Understanding of Paraphrase Types in ChatGPT

### v1

@@ 5 Final Considerations
In this work, we contributed to the understanding of how LLMs can generate variability in linguistic expression through paraphrase types. Using different prompting techniques, we generated 500 paraphrase candidates and asked 15 participants to annotate these paraphrase candidates according to whether the paraphrase type was correctly applied by ChatGPT , the perceived difficulty of the paraphrasing task, and the kind of errors made by ChatGPT at generation. We made the annotations and human preference rankings available through two new datasets (i.e., APTYBaseand APTYRanked).
We found that CoT prompting generally outperforms other methods for generating paraphrases with specific APTs. For paraphrasing tasks perceived as hard for humans, few-shot prompting performed better, while CoT had a remarkable decrease in generation success rate. This suggests that asking ChatGPT to reason about complex paraphrase types (e.g., Synthetic/Analytic Substitution) is ineffective, as the required understanding of these APTs is missing. When ChatGPT made mistakes or added additional changes, they most commonly chose APTs from the morpholexical changes. This is likely because these changes are the most commonly applied for paraphrasing and dominate paraphrasing datasets. We also found that for non-morpholexical changes, ChatGPT rarely confuses changes that can be categorized under the same group (e.g., confusion of one structural change with another instead of a non-structural change). One reason might be that ChatGPT pays increased attention to the properties related to the change so that the retaining of other properties is weighed less. APTYRanked opens an interesting avenue for further research in improving the generation and identification quality of APTs using techniques such as DPO or RHLF. Alternatively, our dataset can be used as a benchmark to develop metrics that better correlate with human preferences.
@@
Although we extend the research in paraphrase types, this study has a few limitations. We consider only a selection of APTs and could not investigate all described types due to the resource-intensive work of human annotation. We focused on the most common ones in the ETPC dataset, which should give a decent proxy for the most relevant types and selected a sample to ensure diversity in the examined types. Additionally, we are limited to one annotation per task for the questionnaire due to effort risking a big variance. However, we used gold examples annotated by each annotator to check the agreement between annotators and to ensure that each annotator performed the annotation attentively. Although we only considered generations from ChatGPT-3.5-turbo-0613, additional LLMs could have been employed. We chose ChatGPT as it is one of the most used models, and this is the first investigation of APT paraphrase generation combined with instruction-trained LLMs. The inclusion of additional models would make the human evaluation too expensive. We do not fine-tune a model on the newly proposed dataset; this paper focuses on the human-evaluated capabilities of an existing model, like ChatGPT.
@@ Appendix A Additional Annotation Information
For annotation phase one, we also used gold questions to evaluate whether the annotators understood the assignment and checked their agreement. The gold examples include completely correct applied paraphrase types, a correctly applied paraphrase with additional applied paraphrase types, a non-paraphrase, and a paraphrase with a different applied APT. The Fleiss Kappa for sense-preservation was 0.92, for correct application 0.68, and for failure reasons 0.61. Determining if a type was applied correctly can depend on details that can be missed, similarly failure reasons are not always trivial to see and depend on the previous answers. There was no annotator whose answers to the gold questions suggested a systematic lack of understanding besides normal human errors, e.g., no annotator answered more than one sense-preserving question incorrectly. For the second annotation phase, i.e., the preferences, Kendall’s coefficient of concordance is 0.52, indicating a moderate agreement on preferences across annotators. Precise preferences are likely highly individual, but trends exist.
@@ B.1 Additional Questions
AQ1. How does the detailed confusion matrix look? Do specific APTs get confused more often than other APTs?
@@
AQ2. Is there a correlation between perceived hardness of a task and confusion?
@@
AQ3. How much do humans and ChatGPT agree on paraphrase type evaluation?

### v2

@@ 5 Final Considerations
In this work, we contributed to the understanding of how LLMs can generate variability in linguistic expression through paraphrase types. Using different prompting techniques, we generated 800 paraphrase candidates. We asked 15 participants to annotate these paraphrase candidates according to whether the paraphrase type was correctly applied by ChatGPT , and the kind of errors made by ChatGPT at generation. We made the annotations and human preference rankings available through two new datasets (i.e., APTYBaseand APTYRanked).
We found that CoT prompting generally outperforms other methods for generating paraphrases with specific APTs. This suggests that asking ChatGPT to reason about complex paraphrase types (e.g., Synthetic/Analytic Substitution) is ineffective, as the required understanding of these APTs is missing. When ChatGPT made mistakes or added additional changes, they most commonly chose APTs from the morpholexical changes. This is likely because these changes are the most common and dominate paraphrasing datasets. We also found that for non-morpholexical changes, ChatGPT rarely confuses changes that can be categorized under the same group (e.g., confusion of one structural change with another instead of a non-structural change). One reason might be that ChatGPT pays increased attention to the properties related to the change so that the retaining of other properties is weighed less. APTYRanked opens an interesting avenue for further research in improving the generation and identification quality of APTs using techniques such as DPO or RHLF. Our experiments with Llama 2 7B-based models show a marked advantage when using human preference data from APTYRankedand DPO and show similar trends to ChatGPT. Alternatively, our dataset can be used as a benchmark to develop metrics that better correlate with human preferences.
@@
Although we extend the research in paraphrase types, this study has a few limitations. We consider only a selection of APTs and could not investigate all described types due to the resource-intensive work of human annotation. We focused on the most common ones in the ETPC dataset, which should give a decent proxy for the most relevant types and selected a sample to ensure diversity in the examined types. Additionally, we are limited to one annotation per task for the questionnaire due to effort risking a big variance. However, we used gold examples annotated by each annotator to check the agreement between annotators and to ensure that each annotator performed the annotation attentively. We focused on generations from ChatGPT-3.5-turbo-0613, as it is one of the most used models, and this is the first investigation of APT paraphrase generation combined with instruction-trained LLMs. We evaluated DPO only for a Llama 7B model; while more models could be interesting, the existing evaluation already shows a marked value of using our preference data. Including additional and larger models would make the human evaluation expensive, so we leave this investigation to future work.
@@ Appendix A Additional Annotation Information
For annotation phase one, we also used gold questions to evaluate whether the annotators understood the assignment and checked their agreement. The gold examples include completely correct applied paraphrase types, a correctly applied paraphrase with additional applied paraphrase types, a non-paraphrase, and a paraphrase with a different applied APT. The Fleiss Kappa for sense-preservation was 0.92, for correct application 0.68, and for failure reasons 0.61. Determining if a type was applied correctly can depend on details that can be missed, similarly failure reasons are not always trivial to see and depend on the previous answers. There was no annotator whose answers to the gold questions suggested a systematic lack of understanding besides normal human errors, e.g., no annotator answered more than one sense-preserving question incorrectly. For the second annotation phase, i.e., the preferences, Kendall’s coefficient of concordance is 0.52, indicating a moderate agreement on preferences across annotators. Precise preferences are likely highly individual, but trends exist. The Fleiss Kappa for correct application is 0.80 for the DPO annotation with two annotators, showing high agreement.
@@ B.1 Additional Questions
AQ1. How does ChatGPT perform on examples perceived as easy or hard by humans?
Ans. The few-shot approach works better in cases where humans evaluate the task as hard, while the CoT-prompted model performs poorly for hard examples
Applying different APTs can be challenging for humans, depending on the precise APT and the base sentence. We explore if the human perceived difficulty of paraphrases also relates to the performance of ChatGPT; that is, are paraphrasing tasks difficult for humans also difficult for ChatGPT?
AQ2. How does the detailed confusion matrix look? Do specific APTs get confused more often than other APTs?
@@
AQ3. Is there a correlation between perceived hardness of a task and confusion?
@@
AQ4. How much do humans and ChatGPT agree on paraphrase type evaluation?